OpenStudy (anonymous):

A polygon has an area of 100 square inches and one of its sides is 6 inches long. If a second similar polygon has an area of 36 square inches, what is the length of the corresponding side in the second polygon?

OpenStudy (anonymous):

i got 7.8 but i dont think its right

OpenStudy (kira_yamato):

Use this formula will do \[\frac{A _{1}}{A_{2}} = \left( \frac{L_1}{L_2} \right)^2\]

OpenStudy (anonymous):

how do i do it.....??? what does "A" stand for?

OpenStudy (kira_yamato):

A = Area L = Length

OpenStudy (anonymous):

im sorry i know i sound real stupid but this stuff is really confusing

OpenStudy (kira_yamato):

Use the attachment...

OpenStudy (anonymous):

thank you so much .... im really not the best at math and everyone always just gives the answer instead of explaing and showing the steps so thank you